# Multi-role Consensus through LLMs Discussions for Vulnerability Detection

### v3

@@ 1 Introduction
As a fundamental process in software engineering, vulnerability detection plays a critical role in ensuring the quality and security of software systems by identifying potential security threats [1, 2]. Nowadays, with the emergence of large language models (LLMs), due to their profound code comprehension, LLMs have also been employed for the application in the realm of vulnerability detection [3, 4].
@@ 4 Conclusion and Future Work
This paper proposes an approach for improving vulnerability detection with LLMs, where LLMs act as different roles in a real-life code review team, discussing the existence and classification of vulnerabilities towards a consensus. Preliminary evaluations have shown notable improvements in recall rates and F1 scores. Future work should integrate in-context learning to guide LLMs to better collaborate in their discussions.

### v4

@@ 1 Introduction
As a fundamental process in software engineering, vulnerability detection plays a critical role in ensuring the quality and security of software systems by identifying potential security threats [1, 2]. Nowadays, with the emergence of large language models (LLMs), due to their profound code comprehension, LLMs have also been employed for the application in the realm of vulnerability detection [3, 4, 5].
@@ 3-A Experiment Settings
Discussion constraints: The maximum discussion depth is set to 5, and the maximum response length is set to 120 tokens.
@@ 4 Conclusion and Future Work
This paper proposes a novel approach for improving vulnerability detection with LLMs, where LLMs act as different roles in a real-life code review team, discussing the existence and classification of vulnerabilities towards a consensus. Preliminary evaluations have shown notable improvements in precision rates, recall rates and F1 scores. Future work should enhance this approach by integrating in-context learning to guide LLMs to better collaborate in their discussions.